# Decomposed Prompting: Unveiling Multilingual Linguistic Structure Knowledge in English-Centric Large Language Models

### v1

@@ 6.1 Multilingual Performance
6.2 English-Centric vs. Multilingual LLMs
| Zero-shot | Few-shot | Avg. | |||
| en | mult. | en | mult. | ||
| LLaMA2-7B | 58.2 | 43.2 | 74.7 | 50.5 | 56.7 |
| BLOOMZ-7B | 20.6 | 17.6 | 44.1 | 36.2 | 29.6 |
| LLaMA2-13B | 59.2 | 48.7 | 65.3 | 48.3 | 55.4 |
| mTk-13B | 47.6 | 43.1 | 57.3 | 44.7 | 48.2 |
Table 4333As an LLM of encoder-decoder structure, mTk is not amenable to direct application of our prob. evaluation designed for decoder-only models; thus, we resort to gen. instead. shows that English-centric LLMs outperform their multilingual counterparts of comparable size by a considerable margin. This superiority, however, is primarily attributed to their proficiency with English linguistic structures knowledge. For languages distant to English and hardly encountered by the English-centric LLMs, such as el,ta,te,yo, BLOOMZ and mTk surpass their English-centric counterparts, with mTk exhibiting enhanced performance across as many as 11 linguistically distant languages. Full results are provided in Appendix C. The observations suggest that multilingual LLMs may possess more robust cross-lingual transferability, but are constrained by their inferior base capabilities.
7 Discussion
Access to LLM Internal Representations
Our ablation study empirically proves that the probability-based evaluation method more accurately reflects the multilingual understanding ability of LLMs than the generation-based method, which merely relies on the output text. However, probability-based evaluation relies on the availability of model output logits, a requirement readily met by open-source LLMs, but not by many other LLMs. This availability of the internal representations of open-source LLMs facilitates intriguing research avenues, for instance, the interpretability of LLM behavior (Saha et al., 2023) and the application of LLMs for Bayesian inference (Li et al., 2023). Hu and Levy (2023) have underscored that direct probability measurement is indispensable in the context of prompting studies. To foster a more transparent and collaborative research environment, better access to the internal workings of LLMs is essential.
LLM’s Path to Multilinguality
This work analyzed the nuances of multilingual performance across different types of LLMs. The exploration of the multilinguality in English-centric LLMs holds significant practical values, particularly when encountering scenarios that demand a model equipped with robust multilingual skills for tasks like reasoning and commonsense understanding. This investigation is pivotal in guiding the decision-making process regarding the foundational model selection for such tasks. The critical question is whether it is more advantageous to commence with an English-centric LLM, which may offer superior understanding abilities in English, and then endeavor to extend these capabilities to additional languages, or to opt for a multilingual LLM that boasts broader language coverage and enhanced multilingual transferability, albeit potentially at the expense of more refined (English) language understanding abilities. Researchers and practitioners should carefully consider the trade-offs between linguistic breadth and depth of language understanding.
@@ 8 Conclusion
In conclusion, our investigation into the multilingual capabilities of English-centric LLMs through the lens of decomposed prompting has yielded significant findings. By systematically dissecting the sequence labeling process into discrete, token-level prompts, we have demonstrated that these models possess a considerable understanding of linguistic structure that extends beyond their predominant English training. Our method outperforms existing iterative prompting techniques in both zero- and few-shot settings, highlighting the efficiency and accuracy of decomposed prompting. The empirical evidence suggests that while English-centric LLMs can effectively engage in multilingual tasks, their performance is nuanced and influenced by the linguistic proximity to English and the design of the prompting strategy. This work not only advances the field of NLP by enhancing our understanding of LLMs’ cross-lingual transfer capabilities but also opens avenues for future research to further improve the inclusivity and adaptability of language models for a diverse range of languages and tasks.
@@
As discussed in the analysis part, our proposed decomposed prompting strategy struggles if the same word occurs twice in a sentence with different POS tags. Besides, the efficiency of decomposed prompting suffers as the length of the input sequence and the complexity of the task increase. Our study uses decomposed prompting methods for part-of-speech (POS) tagging as a means to evaluate the multilingual structural knowledge of English-centric Large Language Models (LLMs). This provides a foundational assessment of the models’ capabilities. Nevertheless, the scope for extending this methodology to probe more intricate aspects of linguistic structure is substantial. Future research could beneficially apply decomposed prompting to the analysis of complex linguistic phenomena, including sentence chunking and syntactic parsing, to gain a deeper understanding of the nuanced capabilities of LLMs in processing and understanding language.
Appendix A Dataset and Languages
A.1 POS Tag Set
@@ A.3 Language Similarity Computation
Appendix B Prompt Details
Zero- and few-shot prompts used in this wor are shown in Figure 8 (decomposed prompting) and Figure 9 (iterative prompting).
Appendix C Full Results
Full experimental results are displayed in Table 6 (Mistral 7B), Table 7 (LLaMA2 7B), Table 8 (LLaMA 13B), Table 9 (BLOOMZ 7B), and Table 10 (mTk 13B).

### v2

@@ 6 Conclusion
In conclusion, we propose decomposed prompting, a simple yet effective prompting method specially designed for sequence labeling tasks, addressing the difficulties of LLM benchmarking on sequence labeling tasks. Our method outperforms iterative prompting techniques in terms of accuracy and efficiency in different experimental settings. By applying decomposed prompting to UDPOS dataset, we probe the multilingual linguistic structure knowledge of English-centric LLMs. Our multilingual investigation reveals that gain from few-shot decomposed prompting is generally more pronounced for languages closer to English.
@@
Although our proposed decomposed prompting method achieves overall remarkable performance in terms of both accuracy and efficiency, it has limitations for some special cases, for example, it can not well handle the case where the same word occurs twice in a sentence with different POS tags. Besides, the efficiency of decomposed prompting suffers as the length of the input sequence and the complexity of the task increase. Our study uses decomposed prompting methods for part-of-speech (POS) tagging as a means to evaluate the multilingual structural knowledge of English-centric Large Language Models (LLMs). This provides a foundational assessment of the models’ capabilities. Nevertheless, extending the application scope of this methodology to probe more intricate aspects of linguistic structure is necessary. Future research could beneficially apply decomposed prompting to the analysis of complex linguistic phenomena, including sentence chunking and named entity recognition, to gain a deeper understanding of the nuanced capabilities of LLMs in processing and understanding language.
Acknowledgements
We want to thank the anonymous reviewers for their valuable feedback. This work was supported by the German Research Foundation (DFG) under grant SCHU 2246/14-1, Munich Center for Machine Learning (MCML), and China Scholarship Council (CSC).
Appendix A Experimental Setup Details
Details of the experimental setup are introduced in this section.
A.1 Dataset and Languages
A.1.1 POS Tag Set
@@ A.1.2 Profile of Languages
A.2 Baselines and Settings
Iterative Prompting (Iter)
Blevins et al. (2023) introduced a structured prompting approach that iteratively labels an entire sentence by appending each predicted label to the context along with the subsequent word. This method is employed as a strong baseline in our study.
Decomposed Prompting (Decom)
To evaluate our proposed approach, we employ the prompt template outlined in §3 to decompose the entire sequence into a set of individual prompts for prediction. In our experiments, we use the 17 POS tags themselves as the label words, i.e., we expect the model to directly predict a tag from the tagset shown in Figure 4 by selecting the tag with the highest logit.
Zero- and Few-Shot Prompting
We devised two experimental scenarios for multilingual prompting—zero-shot and few-shot—to evaluate the performance of both approaches under different conditions. In the zero-shot setting, only an English task instruction is provided alongside the input in the target language. The text in Figure 4, which outlines the tag set information, serves as the instruction in our experiments. In few-shot prompting, we supplement the prompt with a few English demonstrations, structured according to the prompt template of each method. For Decom, we randomly select an example for each tag type from the English training set to create a demonstration. For a fair comparison, the same number of demonstrations are used for the Iter baseline.
@@ A.3 Language Similarity Computation
Appendix B More Details of Decomposed Prompting Method
B.1 Intuition
Figure 6 illustrates the generation of sequence labeling prompts for the German sentence “Viel Erfolg!” via decomposed prompting.
An example of a template function is illustrated as follows.
B.2 Prompt Details
Zero- and few-shot prompts used in this work are shown in Figure 8 (decomposed prompting) and Figure 9 (iterative prompting).
Appendix C Few-Shot Ablation Study
we investigate the impact of the number of few-shot examples on the performance in the decomposed prompting. We randomly select 8 languages (en, de, el, fa, hi, hl, ru, zh) and explore their performance dynamics with the increasing of the few-shot samples. Figure 7 shows that overall, when is small, increasing the number of samples bring performance improvement. As continues to increase, the performance tends to be stable and even gets worse when samples are too many.
Appendix D Full Results
Full experimental results are displayed in Table 4 (Mistral 7B), Table 5 (LLaMA2 7B), Table 6 (LLaMA 13B), Table 7 (BLOOMZ 7B), Table 8 (mTk 13B), and Table 9 (few-shot ablation study).